OpenStudy (anonymous):

Solve: log2^(x+2) = 2 + log2^(x-2)

OpenStudy (anonymous):

I got 3.33

OpenStudy (anonymous):

Is this correct?

OpenStudy (anonymous):

u want a solution or just want to verify ur answer?

OpenStudy (anonymous):

The solution please

OpenStudy (anonymous):

\[\log_{2}( x-2) or \ln 2^{x+1}\]

OpenStudy (anonymous):

THe first one

OpenStudy (anonymous):

But it is a (x+2)

OpenStudy (anonymous):

use this property log a- log = log (a/b)

OpenStudy (kc_kennylau):

no use log(a^b)=b*log(a)

OpenStudy (anonymous):

log2^(x+2) - log2^(x-2) = 2 ?

OpenStudy (anonymous):